A wave with a wavelenght of 54.0 cm is moving at 0.396 m/s. If this wave then increases its wavelenght by 1.05 cm, but maintains the same frequency, what is the new speed of the wave?
1
Expert's answer
2020-10-15T10:41:03-0400

Let's first find the frequency of the wave from the wave speed equation:


"v=f\lambda,"

here, "v" is the speed of the wave, "f" is the frequency of the wave and "\lambda" is the wavelength of the wave.

Then, from this equation we can find the frequency of the wave:


"f=\dfrac{v}{\lambda}=\dfrac{0.396 \ \dfrac{m}{s}}{0.54 \ m} = 0.73 \ Hz."

According to the condition of the question, the wave increases its wavelenght by 1.05 cm, but maintains the same frequency:


"\lambda_{new} = 0.54 \ m + 0.0105 \ m = 0.5505 \ m,""f_{new} = f =0.73 \ Hz."

Finally, we can find the new speed of the wave from the same formula:


"v_{new}=f_{new}\lambda_{new}=0.73\ Hz\cdot 0.5505\ m=0.402 \ \dfrac{m}{s}."

Answer:

"v_{new}=0.402 \ \dfrac{m}{s}."